# Low-cost Embedded Breathing Rate Determination Using 802.15.4z IR-UWB Hardware for Remote Healthcare

### v1

@@ Abstract
Respiratory diseases account for a significant portion of global mortality. Affordable and early detection is an effective way of addressing these ailments. To this end, a low-cost commercial off-the-shelf (COTS), IEEE 802.15.4z standard compliant impulse-radio ultra-wideband (IR-UWB) radar system is exploited to estimate human respiration rates. We propose a convolutional neural network (CNN) to predict breathing rates from ultra-wideband (UWB) channel impulse response (CIR) data, and compare its performance with other rule-based algorithms. The study uses a diverse dataset of 16 individuals, incorporating various real-life environments to evaluate system robustness. Results show that the CNN achieves a mean absolute error (MAE) of 1.73 breaths per minute (BPM) in unseen situations, significantly outperforming rule-based methods (3.40 BPM). By incorporating calibration data from other individuals in the unseen situations, the error is further reduced to 0.84 BPM. In addition, this work evaluates the feasibility of running the pipeline on a low-cost embedded device. Applying 8-bit quantization to both the weights and input/ouput tensors, reduces memory requirements by 67% and inference time by 64% with only a 3% increase in MAE. As a result, we show it is feasible to deploy the algorithm on an nRF52840 system-on-chip (SoC) requiring only 46 KB of memory and operating with an inference time of only 192 ms. Once deployed, the system can last up to 268 days without recharging using a 20 000 mAh battery pack. For breathing monitoring in bed, the sampling rate can be lowered, extending battery life to 313 days, making the solution highly efficient for real-world, low-cost deployments.
@@ 1 Introduction
  • A convolutional neural network (CNN)-based processing pipeline is proposed for breathing determination that shows significant accuracy improvement as compared to basic rule-based processing techniques.
  • The accuracy of the solution is extensively evaluated in various conditions, including an analysis of the generalization performance to unseen people and unknown environments.
  • The impact of model quantization on accuracy, memory consumption, execution time, and energy consumption is modeled and analyzed for embedded devices.
@@ 2 Related work
Year Device Bandwidth (GHz) IEEE comp- liant COTS Embedded impl. Distance Rotation Environ- ment Posture Multiple persons in dataset Multi- person Non static
[8] 2017 NVA6201 2.3 - - - - - - - -
[9] 2019 PulsOn 440 4.3 - - - - - - - -
[10] 2020 X4 NA - - - - - - - -
[11] 2022 X4 NA - - - - - - - -
[12] 2022 X4M03 1.4 - - - - - - - -
[13] 2023 Proprietary 0.5 - - - - - - -
[14] 2020 X4 1.5 - - - - - - -
[6] 2023 NVA6100 NA - - - - - - -
[7] 2024 X4M03 1.4 - - - - - - -
[15] 2024 X4M05 1.4 - - - - - -
[16] 2024 X4M06 NA - - - - - -
[17] 2022 X4m200 1.4 - - - - - -
[18] 2017 NVA6201 2.3 - - - - - -
@@
Table 1: Overview of related work, categorized by hardware type and application context. This work is the first to provide a truly low-cost, IEEE 802.15.4z-compliant, off-the-shelf available solution that can run on embedded hardware while coping with a variety of real-life environments and conditions.
@@
Most prior work utilizes algorithm-based breathing rate estimators. A notable exception is Zhou et al. [19], which addresses a comparable number of challenges compared to this paper but lacks evaluations in multiple environments or with subjects in different postures. However, it does excel in analyzing non-static scenarios and multi-person setups. To tackle these challenging situations, the paper proposes an Attention-Based CNN-long short-term memory (LSTM) network to filter out human motion, after which a rule-based approach is used to obtain the BR. The computational complexity and low-cost aspects are not discussed, as the Novelda hardware and a very complex neural network (NN) is used.
The work described in this paper stands out from all works in Table 1 by using COTS IEEE 802.15.4z-compliant hardware, validating the system with multiple real-life considerations in mind and also analyzing the processing requirements of the solution. Compared to prior work, this study is the first to focus on low-cost and embedded hardware compatibility aspects.
@@ 3 System model
This study employs a radar-based system with a transmitter emitting signals that reflect off objects or people in the room. These reflections will be influenced by respiratory movements of the person in front of the radar and are subsequently captured by the receiver. The varying distance between the radar and the chest of the target is then used to extract the breathing rate [21]. Equation 1 shows mathematically how this distance is modelled.
@@ 4 Measurement campaign and dataset description
For evaluating the different algorithms, data was collected using the Qorvo DW3000, specifically the QM33120WDK1 development kit (DK), acting as a representative low-cost, COTS hardware device. The setup includes separate transmitter and receiver devices, each equipped with a Nordic nRF52840 DK and a DW3000 transceiver shield. The transmitter uses an omnidirectional antenna, while the receiver employs a directional antenna to focus on reflections coming from the desired direction. The system operates at a center frequency of 6.489 GHz with a bandwidth of 500 MHz. The slow-time sampling frequency is 77.5 Hz, and the receiver captures IQ samples at 1 ns intervals. Each fast-time sample represents 30 cm in distance, with each CIR containing 41 fast-time samples and the first peak always aligned at index 3. Samples before the first peak can be considered noise. The maximum distance the system can reach is equal to: $((41-4)*30~{}cm)/2=555~{}cm$. A Plux respiratory effort belt [22] serves as ground truth, measuring the accurate BR through belt stretch.
@@ 5.1 Preprocessing
The preprocessing pipeline is shown in Figs. 2a to 2e. This pipeline processes the samples captured as IQ samples. For each slow-time sample, 41 fast-time samples are taken, forming a single CIR. Once collected, the amplitude of these samples is calculated. The resulting amplitude versus fast-time is shown in Fig. 2a. While phase information is also available, it is discarded due to unsynchronized transmitter and receiver clocks. These CIR s are then combined into a matrix, of which a heatmap representation is shown in Fig. 2b. Each row in the matrix corresponds to a CIR (the “slow time”’), and each column captures variations across time at a certain distance (the “fast time”). The maximum value of the slow-time index of the matrix is 120 seconds, representing all CIR samples collected from a single person in a specific environment and setup. To calculate the breathing rate over a shorter duration than the full experiment, a 30-second windowing operation is applied to the CIR matrix with a stride of 15 seconds. This window size was empirically determined to be an optimal balance between the dataset size required for model training and the accuracy of the fast fourier transform (FFT) that will be applied later in the pipeline. Using an overlapping windowing approach results in more training samples, which is advantageous for training a NN. Careful splitting between train and validation samples in our validation strategy will ensure no data leakage is introduced, which is further detailed in Section 6. An example of a window is shown in Fig. 2c. In this representation, it is clear that a periodic signal is present around range bin 8. In the next step, an FFT is applied to each column, generating a frequency spectrum for each distance bin. The obtained representation is shown on Fig. 2d, which is referred to as a range-frequency map. Peaks at the target breathing frequency are expected at the range index where a person is present. To focus on breathing rate-specific frequencies, signals below 0.09 Hz or above 0.5 Hz should be filtered out. When zooming in to this frequency range, peaks are indeed present in the example. Since the first peak index is present at index 3, the first 4 range bins can also be removed. Both the frequency filtering and range bin removal can be implemented with an image crop operation. The result of this operation is shown in Fig. 2e, showing which movement frequencies are present at each distance. The final step involves applying Min-Max normalization to scale all values between 0 and 1. This ensures consistent data scales across all CIR windows, which will benefit NN approaches. The min and max values are derived solely from the training data to prevent data leakage. The normalized CIR window then serves as input to the breathing detection algorithms, which in our case will be either peak finding techniques or neural networks.
@@ 5.3 Convolutional neural network based machine learning
The final method evaluated is a CNN. Among existing studies, only Zhou et al. [19] proposes an attention-based CNN-LSTM solution evaluated on a similar amount of real-world considerations as our dataset. A more simple CNN was chosen in this paper due to its significantly lower processing requirements, aligning better with the low-cost aspect of this paper. The model architecture and optimized parameters are detailed in Table 2. Various configurations of parameter sizes, kernel shapes, and layer counts were tested, with the best-performing design summarized in the table. A grid search was used to optimize normalization and regularization techniques, including batch normalization, dropout, and kernel regularization. The best-performing settings are also provided in Table 2. To ensure the robustness of the results, each model was trained 10 times over. The model that scored the median validation loss was selected. All subsequent reported results follow this training strategy.
@@ 6.1 Generalization to new persons
The first analysis evaluates the leave-one-person-out cross-validation strategy. In this approach, 1 participant is excluded from the training set in each fold and used for validation. This process is repeated for all 16 participants, producing 1 model per participant. The results are summarized in Table 3, showing the mean L1 errors and their standard deviation. For the highest peak method, mean L1 errors range from 3.10 BPM to 5.58 BPM. The accumulated distance-frequency map approach reduces mean L1 errors slightly, ranging from 2.75 BPM to 5.32 BPM. The weighted average algorithm shows L1 errors from 3.29 BPM to 4.46 BPM. Among the rule-based methods, selecting the highest peak from the accumulated frequency bins proves most effective. The table also includes results for our CNN model, where mean L1 errors range from 0.97 BPM to 2.51 BPM. This demonstrates a significant improvement compared to the rule-based solutions. Fig. 3 illustrates in which conditions the CNN outperforms traditional rule-based approaches. Fig. 3a shows a range-frequency representation in ‘simple’ conditions, where minimal noise was present. Hence, the breathing frequency peak is clearly visible and can be identified by most algorithms. However, in more challenging scenarios, where the noise becomes more pronounced, some noise frequency components may exceed the breathing signal. One example is shown in Fig. 3b. In these cases, the CNN performs better by not only analyzing the highest peak, but it can also take the higher harmonic components into account.
When compared to prior work, Farsaei et al. [13] and Numan et al. [20] report mean L1 errors of $<$BPM and 1.22$\pm$0.63 BPM, respectively. In comparison, the rule-based approaches evaluated here show worse L1 scores. Our CNN comes closer to this error but still results in a higher L1 error. This difference can be attributed to the inclusion of significantly more challenging scenarios and greater variability in BR s in our study. A fairer comparison is with the results from Zhou et al. [19], who report a mean L1 error of 2.3$\pm$1.9 BPM across various challenging scenarios. Our proposed CNN outperforms their results, with all validation folds in Table 3 achieving lower mean L1 scores.
Highest peak Accumulated highest peak Accumulated weighted average CNN
Pers. mean (BPM) std (BPM) mean (BPM) std (BPM) mean (BPM) std (BPM) mean (BPM) std (BPM)
1 3.94 4.20 3.13 2.79 3.89 2.18 1.65 1.24
2 4.20 4.02 4.05 3.85 4.34 2.51 2.08 1.52
3 5.38 4.34 5.32 4.89 3.54 2.50 2.51 2.33
4 4.31 4.14 2.87 3.32 3.85 2.27 1.51 1.30
5 4.57 4.25 3.76 3.97 3.57 2.39 1.59 1.27
6 4.15 3.48 3.83 3.54 3.64 2.63 2.06 1.5
7 3.61 3.15 2.97 3.18 3.57 1.88 0.97 0.83
8 4.94 4.19 3.58 3.70 3.80 2.66 2.17 1.57
9 5.58 4.67 4.10 3.67 3.99 2.56 2.05 1.80
10 4.19 4.53 2.75 3.06 4.26 2.32 1.35 1.18
11 3.10 3.40 3.19 3.65 4.25 2.34 1.64 1.48
12 3.67 3.14 3.33 2.99 4.46 2.19 1.83 1.34
13 4.19 4.49 2.98 2.89 4.10 2.27 1.89 1.29
14 3.29 3.26 2.80 3.10 3.98 2.21 1.54 1.38
15 3.52 3.78 2.86 3.41 3.29 2.42 1.34 1.11
16 3.61 3.74 3.01 3.92 4.22 2.07 1.19 0.97
@@
Table 3: Mean and standard deviation of the L1 error obtained when predicting the breathing rate of unseen persons. The CNN performs better than rule-based techniques, with mean accuracies between 0.97 and 2.51 BPM and lower standard deviations.

### v2

@@ Abstract
Respiratory diseases account for a significant portion of global mortality. Affordable and early detection is an effective way of addressing these ailments. To this end, a low-cost commercial off-the-shelf (COTS), IEEE 802.15.4z standard compliant impulse-radio ultra-wideband (IR-UWB) radar system is used to estimate human respiration rates. We propose a convolutional neural network (CNN) specifically adapted to predict breathing rates from ultra-wideband (UWB) channel impulse response (CIR) data, and compare its performance with both other rule-based algorithms and model-based solutions. The study uses a diverse dataset, incorporating various real-life environments to evaluate system robustness. To facilitate future research, this dataset will be released as open source. Results show that the CNN achieves a mean absolute error (MAE) of 1.73 breaths per minute (BPM) in unseen situations, significantly outperforming rule-based methods (3.40 BPM). By incorporating calibration data from other individuals in the unseen situations, the error is further reduced to 0.84 BPM. In addition, this work evaluates the feasibility of running the pipeline on a low-cost embedded device. Applying 8-bit quantization to both the weights and input/ouput tensors, reduces memory requirements by 67% and inference time by 64% with only a 3% increase in MAE. As a result, we show it is feasible to deploy the algorithm on an nRF52840 system-on-chip (SoC) requiring only 46 KB of memory and operating with an inference time of only 192 ms. Once deployed, an analytical energy model estimates that the system, while continuously monitoring the room, can operate for up to 268 days without recharging when powered by a 20 000 mAh battery pack. For breathing monitoring in bed, the sampling rate can be lowered, extending battery life to 313 days, making the solution highly efficient for real-world, low-cost deployments.
@@ 1 Introduction
  • A convolutional neural network (CNN)-based processing pipeline is proposed for breathing determination that shows significant accuracy improvement as compared to rule-based, and model-based processing techniques.
  • The proposed solution is evaluated with an improved strategy considering multiple environments, subjects, and orientations, providing a more thorough assessment of generalization to unseen conditions.
  • The algorithmic complexity of each processing step is analyzed, providing insights into scalability and computational feasibility on resource-constrained devices.
  • The effects of model quantization on accuracy, memory usage, execution time, and energy consumption are modeled and evaluated, highlighting practical trade-offs for embedded deployment.
  • An analytical energy model is developed, covering the radar hardware, signal-processing pipeline and the model inference stage, enabling accurate predictions of execution time, memory footprint, and energy consumption.
@@ 2 Related work
Year Device
Bandwidth
IEEE
comp-
liant
COTS
Embedded
impl.
Dist-
ance
Rotat-
ion
Environ-
ment
Post-
ure
Multiple
persons
in dataset
Multi-
person
Non
static
AI-
based
[8] 2017 NVA6201 2.3 - - - - - - - - -
[9] 2019 PulsOn 440 4.3 - - - - - - - - -
[10] 2020 X4 NA - - - - - - - - -
[11] 2022 X4 NA - - - - - - - - -
[12] 2022 X4M03 1.4 - - - - - - - - -
[13] 2023 Proprietary 0.5 - - - - - - - -
[14] 2020 X4 1.5 - - - - - - - -
[6] 2023 NVA6100 NA - - - - - - - -
[7] 2024 X4M03 1.4 - - - - - - - -
[15] 2024 X4M05 1.4 - - - - - - -
[16] 2024 X4M06 NA - - - - - - -
[17] 2022 X4m200 1.4 - - - - - - -
[18] 2017 NVA6201 2.3 - - - - - - -
@@
Table 1: Overview of related work, categorized by hardware type and application context. This work is the first to provide a truly low-cost, IEEE 802.15.4z-compliant, off-the-shelf available solution that can run on embedded hardware while coping with a variety of real-life environments and conditions. Indicated by the final column, using an AI-approach in this context has not been explored extensively.
@@
The final column of Table 1 shows whether a solution uses AI. All but one prior work utilizes a model-based BR estimator. This exception is Zhou et al. [19], which addresses a comparable number of challenges compared to this paper but lacks evaluations in multiple environments or with subjects in different postures. However, it does excel in analyzing non-static scenarios and multi-person setups. Their approach uses an Attention-Based CNN-long short-term memory (LSTM) model to filter out motion. Afterwards, the filtered data is analyzed with a rule-based BR estimator. In contrast to our work the authors do not analyze the computational costs or provide energy models to evaluate the suitability for embedded implementations. The lack of environmental testing also raises major concerns about generalizability, a key hurdle in AI-driven solutions.
The work described in this paper stands out from all works in Table 1 by using COTS IEEE 802.15.4z-compliant hardware, validating the system with multiple real-life considerations in mind and also analyzing the processing requirements of the solution. Compared to prior work, this study is the first to focus on analyzing the computational and energy requirement aspects required for embedded hardware compatibility.
@@ 3 System model
This study employs a radar-based system with a transmitter emitting signals that reflect off objects or people in the room. These reflections will be influenced by respiratory movements of the person in front of the radar and are subsequently captured by the receiver. The varying distance between the radar and the chest of the target is then used to extract the BR [21]. Equation 1 shows mathematically how this distance is modelled.
@@ 4 Measurement campaign and dataset description
For evaluating the different algorithms, data was collected using the Qorvo DW3000, specifically the QM33120WDK1 development kit (DK), acting as a representative low-cost, COTS hardware device. The setup includes separate transmitter and receiver devices, each equipped with a Nordic nRF52840 DK and a DW3000 transceiver shield. The transmitter uses an omnidirectional antenna, while the receiver employs a directional antenna to focus on reflections coming from the desired direction. The system operates at a center frequency of 6.489 GHz with a bandwidth of 500 MHz. The slow-time sampling frequency is 77.5 Hz, and the receiver captures IQ samples at 1 ns intervals. Each fast-time sample represents 30 cm in distance, with each CIR containing 41 fast-time samples and the first peak always aligned at index 3. Samples before the first peak can be considered noise. The maximum distance the system can reach is equal to: ((414)30cm)/2=555cm((41-4)*30~cm)/2=555~cm. A Plux respiratory effort belt [22] serves as ground truth, measuring the accurate BR through belt stretch.
@@ 5.1 Preprocessing
The preprocessing pipeline is shown in Figs. 2a to 2e. This pipeline processes the samples captured as IQ samples. For each slow-time sample, 41 fast-time samples are taken, forming a single CIR. Once collected, the amplitude of these samples is calculated. The resulting amplitude versus fast-time is shown in Fig. 2a. While phase information is also available, it is discarded due to unsynchronized transmitter and receiver clocks. These CIR s are then combined into a matrix, of which a heatmap representation is shown in Fig. 2b. Each row in the matrix corresponds to a CIR (the “slow time”’), and each column captures variations across time at a certain distance (the “fast time”). The maximum value of the slow-time index of the matrix is 120 seconds, representing all CIR samples collected from a single person in a specific environment and setup. To calculate the breathing rate over a shorter duration than the full experiment, a 30-second windowing operation is applied to the CIR matrix with a stride of 15 seconds. This window size was empirically determined to be an optimal balance between the dataset size required for model training and the accuracy of the fast fourier transform (FFT) that will be applied later in the pipeline. Using an overlapping windowing approach results in more training samples, which is advantageous for training a neural network (NN). Careful splitting between train and validation samples in our validation strategy will ensure no data leakage is introduced, which is further detailed in Section 6. An example of a window is shown in Fig. 2c. In this representation, it is clear that a periodic signal is present around range bin 8. In the next step, an FFT is applied to each column, generating a frequency spectrum for each distance bin. The obtained representation is shown on Fig. 2d, which is referred to as a range-frequency map. Peaks at the target breathing frequency are expected at the range index where a person is present. To focus on breathing rate-specific frequencies, signals below 0.09 Hz or above 0.5 Hz should be filtered out. This band-pass operation also suppresses low-frequency components arising from static objects and slow-varying factors such as temperature or background changes. When zooming in to this frequency range, peaks are indeed present in the example. Since the first peak index is present at index 3, the first 4 range bins can also be removed. Both the frequency filtering and range bin removal can be implemented with an image crop operation. The result of this operation is shown in Fig. 2e, showing which movement frequencies are present at each distance. The final step involves applying Min-Max normalization to scale all values between 0 and 1. This ensures consistent data scales across all CIR windows, which will benefit NN approaches. The min and max values are derived solely from the training data to prevent data leakage. The normalized CIR window then serves as input to the breathing detection algorithms, which in our case will be either rule-based techniques or neural networks.
@@ 5.3 Convolutional neural network based machine learning
Among existing studies, only Zhou et al. [19] proposes an attention-based CNN-LSTM solution evaluated on a similar amount of real-world considerations as our dataset. A more simple CNN was chosen in this paper due to its significantly lower processing requirements, aligning better with the low-cost aspect of this paper. The model architecture and optimized parameters are detailed in Table 2. Various configurations of parameter sizes, kernel shapes, and layer counts were tested, with the best-performing design summarized in the table. A grid search was used to optimize normalization and regularization techniques, including batch normalization, dropout, and kernel regularization. The best-performing settings are also provided in Table 2. To ensure the robustness of the results, each model was trained 10 times over. The model that scored the median validation loss was selected. All subsequent reported results follow this training strategy.
Although not previously applied to IR-UWB-based BR estimation, several resource-efficient model-based approaches exist, such as Random Forests and lightweight CNN s like MobileNet. To provide constrained-AI baselines, we applied both methods to our dataset. The Random Forest was implemented on flattened input features, while MobileNet used pretrained convolutional layers (ImageNet) with adapted fully connected layers for regression. To provide an additional CNN baseline without an explicit emphasis on deployability, we also evaluated a ResNet architecture with pretrained convolutional layers (ImageNet) and modified dense layers. A key comparison factor is model size: our proposed CNN requires only 141 KB, whereas the MobileNet and ResNet implementations used in this study occupy 23.7 MB and 89.6 MB, respectively. Beyond compactness, our model is specifically tailored to the input size and characteristics of IR-UWB radar signals. In particular, its filters are designed to effectively capture frequency components across different range bins in the range–frequency maps. Its compactness also lowers the risk of overfitting, thereby supporting better generalization across diverse environments. More complex architectures (e.g., LSTM or transformers) were not considered, as they are incompatible with the embedded, resource-limited focus of this work.
@@ 6.1 Generalization to new persons
The first analysis evaluates the leave-one-person-out cross-validation strategy. In this approach, 1 participant is excluded from the training set in each fold and used for validation. This process is repeated for all 16 participants, producing 1 model per participant. The results are summarized in Table 3, showing the mean L1 errors and their standard deviation. For the highest peak method, mean L1 errors range from 3.10 BPM to 5.58 BPM. The accumulated distance-frequency map approach reduces mean L1 errors slightly, ranging from 2.75 BPM to 5.32 BPM. The weighted average algorithm shows L1 errors from 3.29 BPM to 4.46 BPM. Among the rule-based methods, selecting the highest peak from the accumulated frequency bins proves most effective.
The table also includes results from other AI-based approaches previously mentioned in Section 5. The results in the rightmost column of Table 3 show the performance of our proposed model. Overall, model-based solutions clearly outperform rule-based methods. Fig. 3 highlights conditions where the model-based solutions surpasses traditional approaches. In ‘simple’ conditions with little noise (Fig. 3a), the breathing frequency peak is easily identified by most algorithms. In more challenging scenarios, however, noise components may dominate the breathing signal (Fig. 3b). In these cases, the model-based solutions perform better by not only analyzing the highest peak, but also taking the higher harmonic components into account. Among all investigated solutions, our proposed CNN achieves the best performance, with mean L1 errors between 0.97 BPM and 2.51 BPM, demonstrating a clear improvement over both rule-based methods and current AI solutions. This clearly demonstrates that the architecture proposed in this work is well-optimized for effective performance on IR-UWB data.
[…]
Collecting datasets in healthcare settings is resource-intensive and requires sufficient participant diversity to ensure statistical validity and generalizability. To assess how performance scales with sample size, we employed an extended leave-k-people-out cross-validation with multiple permutations of training–testing splits. Results show that L1 error consistently decreases as the number of training participants increases, though with diminishing returns: expanding from 1 to 5 participants reduced error by 0.262 BPM, from 6 to 10 by 0.261 BPM, and from 10 to 16 by 0.154 BPM. This indicates that while additional data may enhance performance, the incremental gains are expected to diminish quickly as more individuals are added.
Rule based AI based Constrained AI based
Highest
peak
Accumulated
highest
peak
Accumulated
weighted
average
ResNet50 Random forest Mobilenet
CNN
(Our solution)
Person
mean
std
mean
std
mean
std
mean
std
mean
std
mean
std
mean
std
1 3.94 4.20 3.13 2.79 3.89 2.18 2.04 1.51 2.05 1.30 2.03 1.49 1.65 1.24
2 4.20 4.02 4.05 3.85 4.34 2.51 2.59 1.6 2.40 1.43 2.55 1.72 2.08 1.52
3 5.38 4.34 5.32 4.89 3.54 2.50 3.09 2.2 2.82 2.31 2.77 2.1 2.51 2.33
4 4.31 4.14 2.87 3.32 3.85 2.27 2.47 1.28 2.13 1.41 2.47 1.29 1.51 1.30
5 4.57 4.25 3.76 3.97 3.57 2.39 2.7 1.56 2.34 1.36 2.59 1.56 1.59 1.27
6 4.15 3.48 3.83 3.54 3.64 2.63 2.74 1.21 2.57 1.32 2.69 1.63 2.06 1.5
7 3.61 3.15 2.97 3.18 3.57 1.88 1.97 1.19 1.68 1.19 1.96 1.24 0.97 0.83
8 4.94 4.19 3.58 3.70 3.80 2.66 2.88 1.4 2.46 1.48 2.87 1.37 2.17 1.57
9 5.58 4.67 4.10 3.67 3.99 2.56 2.57 1.4 2.47 1.52 2.58 1.39 2.05 1.80
10 4.19 4.53 2.75 3.06 4.26 2.32 2.23 1.38 1.88 1.22 2.18 1.39 1.35 1.18
11 3.10 3.40 3.19 3.65 4.25 2.34 2.42 1.6 2.16 1.30 2.34 1.75 1.64 1.48
12 3.67 3.14 3.33 2.99 4.46 2.19 2.3 1.64 2.12 1.26 2.28 1.3 1.83 1.34
13 4.19 4.49 2.98 2.89 4.10 2.27 2.02 1.32 1.94 1.29 2.01 1.51 1.89 1.29
14 3.29 3.26 2.80 3.10 3.98 2.21 2.27 1.44 2.20 1.52 2.28 1.42 1.54 1.38
15 3.52 3.78 2.86 3.41 3.29 2.42 2.56 1.45 1.99 1.30 2.52 1.52 1.34 1.11
16 3.61 3.74 3.01 3.92 4.22 2.07 2.32 1.96 2.24 1.20 2.29 1.68 1.19 0.97
Average 4.14 3.92 3.41 3.50 3.92 2.34 2.45 1.51 2.22 1.40 2.40 1.52 1.71 1.38
@@
Table 3: Mean and standard deviation of the L1 error obtained when predicting the breathing rate of unseen persons. The CNN performs better than rule-based and other AI-based techniques, with mean accuracies between 0.97 and 2.51 BPM and lower standard deviations.